Opened 9 years ago

Last modified 4 years ago

#402 assigned flight processing

CAM11_03, flight day 141/2011, Madrid

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/spain_portugal/CAM11_03-2011_141_Madrid

Data arrived from ARSF via network transfer on 24/05/2011

CAM11_03 Scientific objective: We wish to measure how these forests change over gradients of topography‐controlled climate and soil properties.

CAM11_03 Priority: Alpha 4h

PI: David Coomes

Sensors:

Camera (24/06/2011)
Eagle (27/06/2011)
Hawk (27/06/2011)
LiDAR (15/08/2011)

Note: This project is actually part of EU11_03 but has a different project code as it was funded separately, and has higher priority.

Attachments (1)

CAM11_03_141.jpg (94.6 KB) - added by emca 8 years ago.
scan_angle_issue

Download all attachments as: .zip

Change History (26)

comment:1 Changed 9 years ago by knpa

  • Description modified (diff)

comment:2 Changed 9 years ago by iopa

Taking LiDAR and camera.

comment:3 Changed 8 years ago by iopa

Camera is ready for checking @ ~arsf/arsf_data/2011/flight_data/spain_portugal/CAM11_03-2011_141_Madrid/processing/delivery/camera-CAM11_03-100-20110615

Raw images taken during preparation and removed ;
1-28
Raw images with cloud or saturation and removed ;
155,194,276,607-615,699

comment:4 Changed 8 years ago by knpa

  • Milestone set to 2011 data processing completion
  • Owner set to iopa
  • Status changed from new to assigned

Taking Camera delivery check.

comment:5 Changed 8 years ago by knpa

Delivery check complete. All looks good apart from:

  1. Looks like the calibration certificate from last year. Is that OK?
  2. Noticed in the tiff tags that the gps time is given as 10/04/11, not 10/05/11 when the project was flown.

comment:6 Changed 8 years ago by benj

LiDAR processing currently blocked until we get a new boresight flight

comment:7 Changed 8 years ago by anch

Hyperspectral ready for delivery check

comment:8 Changed 8 years ago by knpa

Taking hyperspectral delivery check.

comment:9 Changed 8 years ago by knpa

Delivery check complete.

All looks good apart from:

  1. Logsheet top section has lots missing. Can fill this out from the scanned logsheet.
  2. Readme: aplcorr command needs altering. e141011b_post_processed_nav.bil should be e141011b_nav_post_processed.bil (script needs editing also)
  3. Blank bands at 127 and 0 (eagle)
  4. Blank band at 0 (hawk)

comment:10 Changed 8 years ago by anch

Issues with delivery resolved.

Now putting the delivery on a disk

comment:11 Changed 8 years ago by anch

  • Description modified (diff)

Hawk and Eagle data sent to PI on 27/06/2011

comment:12 Changed 8 years ago by iopa

  • Description modified (diff)

comment:13 Changed 8 years ago by emca

Starting LiDAR processing. Navigation has already been processed.

comment:14 Changed 8 years ago by emca

LiDAR lines 7 and 8 (lines '3-4' and '1-2' on the logsheet) have short sections where there was an issue with scan angle.

The scan angle should oscillate regularly between approx. +-20 degrees. During flight line 8 the scan angle reduces down to 1, then goes up to ~34/35 before stabilising again. A similar pattern occurs several times over the two lines (the attached image explains better).

Both of these lines were repeat lines so the small sections in question are covered by other lines.

Changed 8 years ago by emca

scan_angle_issue

comment:15 Changed 8 years ago by emca

CORRECTION: The scan line issue referred to above occurs on lines 6 and 7 - these are lines '5-6b' and '3-4b' on the logsheet.

LiDAR data is now ready for delivery check.

comment:16 Changed 8 years ago by jaho

LIDAR delivery check:

In addition to mentioned scan angle issue in lines 6 and 7, there seem to be some gaps and minor artifacts in lines: 3, 4 and 5.

Other than that the data looks good to go.

comment:17 Changed 8 years ago by emca

Data gaps in lines 3,4 5 are due to cloud cover. Have added this info to readme.

Data is on disk 112 and will be sent along with data from day 142.

comment:18 Changed 8 years ago by emca

Above comments refer to LiDAR data delivery

comment:19 Changed 8 years ago by emca

  • Description modified (diff)

Delivered to PI 15/08/2011 along with data from day 142 (ticket #401)

comment:20 Changed 8 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:21 Changed 8 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.055-0.095
2-0.035-0.045
3-0.025-0.055
4-0.055-0.055
5-0.055-0.055
6-0.055-0.055
7-0.045-0.045
8-0.055-0.085
9-0.025-0.055
10-0.025-0.055
11-0.055-0.055

comment:22 Changed 8 years ago by knpa

Beginning archiving.

comment:23 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from iopa to knpa

comment:24 Changed 8 years ago by knpa

Archiving complete. Made available to NEODC.

comment:25 Changed 4 years ago by dac

There was an identical copy of this project under /users/rsg/arsf/arsf_data/2011/flight_data/uk - have removed.

Note: See TracTickets for help on using tickets.